Reports 1-1 of 1 Clear search Modify search
MIF (ITF Control)
masayuki.nakano - 10:49 Friday 14 February 2020 (12961) Print this report
CARM loop cleaned up

[Akutsu, Nakano]

New record!!

The CARM common mode servo gain configuration was not optimized and it has 10 times more servo noise. That limited the sensitivity and the binary range was improved up to 390 kpc.

  • Before the input gain was -20 dB (20 dB attenuator between the CMS and the IQ demod) and the output gain was 30 dB. This means we had 10 times larger servo noise. So, I increased input gain to 0 dB (removing attenuator) and reduce the output gain to 10 dB.
  • I put two HPF reported by Akutsu san in klog12959. However, I asked Yamat-sensei to remove one of them due to the phase delay.
  • Still, we have unknown low frequency noise at I/Q demod board. We will continue to investigate it.
  • The lock stability is amazingly good even without ASC. It is an advantage of underground, I guess.


Comments to this report:
takaaki.yokozawa - 11:19 Friday 14 February 2020 (12962) Print this report

Raipd analysis of the Bruco,

Around the 100Hz, BS TM damp would be one of the main noise source.

Images attached to this comment
takahiro.yamamoto - 7:51 Monday 17 February 2020 (12994) Print this report

Estimated sensitivity had been worse as a factor of 1.5 because the exchange of the coil driver chassis had been missed in the calibration.
(Super high power coil driver had been used for ETMY TM.)

Calibration and online h(t) reconstruction was fixed as I reported on klog12992.
Unfortunately, IFO lock is difficult today due to the microseismic disturbance, but correct BNS range can be estimated as roughly 600kpc.

Search Help